CRY+VIS Interface Meeting
Date: 2022/2/4 13:00-13:35
Participants: Takahashi, Uchiyama, Node, Hirata, Yasui, Sato, Tomaru, Yokozawa, Ikeda, Kimura, Suzuki, Aso, Ushiba
- IX: Glitch in the F3 signals. Windshield for the MN oplev. Cooler is ready.
- EX: Health check will be done after the vacuum issue. Repair of the feedthrough is scheduled on 2/9-10.
- IY: Height adjustment was done on 2/1. Height measurement was scheduled in the next week.
- EY: Height measurement was done on 2/1. MN oplev was set. Preparation for the accelerometers.
- Replacement of flip cables: Type-C. Cables were delivered.
- TM height setting: Measurement for ITMY.
- Oplev: Setup of the TM Oplev in IY after the TM alignment in late Feb.
- ACC: Preparation for EY.
- Commissioning in air: Return of GRY. Fluctuation of ETMY is large. MN oplev control is necessary.
- Commissioning in vacuum:
- Overall commissioning:
- Acceptance check:
- PF actuator: At the in-vacuum commissioning.
- New vacuum pumps for V-chambers: EXV (transportation on 2/7, implementation on 2/8-11)
- Heater control: Test in IXV/EXV.
- Power source for Oplev: IY.
- TMS: This will be rescheduled due to the lost parts. Leak test should be done beforehand.
Date: 2022/2/4 14:00-15:00
Participants: Takahashi, Uchiyama, Ushiba, Ishizaki, Fabian, Aso, Tanaka, Sato, Hirata
- Adjusted the height of ITMY and F1 damper.
- Investigating the noise source in the accelerometers.
- Repaired the F0 yaw FR in EY.
- BS F1 LVDT:
- After calibration last week, I updated the medm values accordingly, however, the amount of actuation needed to move the keystone to the setpoint was too high (-26,000 counts).
- In order to solve this, I modified the LVDT calibration offset to yield zero, when the keystone is at the nominal 3D-CAD position. I achieved this using the old and new calibration data.
- The difference between the setpoints is only 324 um, which is negligible in terms of the uncertainties we have when we measure the height of the optic.
- The amount of actuation reduced to around -8,000 counts, and will decrease more when the system goes into vacuum.
- Terrence reported there is some residual motion in the BS suspension. The oplev measures a 50 mHz yaw oscillation, which the damper ring is supposed to damp.
- I assessed the performance of the damper ring with the F0 keystone at two different heights, and from visual inspection of the time series, it seems to be working.
- The residual motion is even there with the control system off.
- Currently, we have the impression it might be an environmental influence. The chamber is not properly shielded against air currents.
- We need to repeat some measurements when there is no one working in the tunnel, and if this does not produce good results, we would like to close the chamber with the clean clothes rather than with the improvised aluminium duct.
We inspected PR2 traverser, checked a cable and adjusted the height of PR3 suspended breadboard.
- Type-A team: Accelerometers in IY/EY.
- Type-B team: Cable for BF FR in BS. Work on the vacuum pump by the SR chamber on 2/7 or 8.
(Date in Kamioka)
The week from 2/7
- Takahashi: 2/8-10
- Sato: 2/8-10
- Hirata: 2/8-10
The week from 2/14
- Takahashi: 2/15-18
- Sato: 2/15-18
- Hirata: 2/15-18